An athlete runs 105 m south and 80 m west. How far is he from the starting point?
Answer by mananth(16946)

An athlete runs 105 m south and 80 m west. How far is he from the starting point?
after running 105 m he turns 90 deg. and runs 80 m.
80 m and 105 m for the legs of a right triangle.
the distance from starting point is the hypotenuse.
105 ^2 +80^2 = hyp ^2
11025 +6400= hyp^2
17425=hyp^2
hyp = 132 m the distance from the starting point